If I have an initial concentration of NaOH of 1.25M in 15mL, and add 10mL water to get a final concentration of 25mL, what's the final concentration of the NaOH?

I did C1V1 = C2V2 and got 0.5M

Is that correct?
Answer #1

Here, C1 = 1.25 M , C2 = 0 M, V1 = 15 mL, V2 = 10 mL.

Then volume of mixture formed is:

Vmix = 15 + 10 = 25 mL

Calculate the concentration of the mixture formed as follows:

CmixVmix = C1V1 + C2V2

Cmix*25 = 1.25*15 + 0*10

Cmix = (1.25*3)/5 = 0.25*3 = 0.75 M

Therefore, the concentration of mixture formed is 0.75 M. Answer is 0.75 M.

Please give good rating.